# Shenzhen, China

### v5

@@ 0.B.1 Evaluation metric implementation
As CLIP has been designed to align text and image modalities, we also leverage its text encoder for EEG classification using the text embeddings of categories. This approach utilizes CLIP’s text encoding capabilities to facilitate EEG classification. We conducted zero-shot classification tests on the THINGS-EEG dataset. We employed Top-K accuracy as a metric for performance evaluation. Specifically, we assessed performance based on the Top-k (where k=1, 5) predictions. This means a classification is considered correct if the true category is among the model’s Top-k predicted categories. We conducted tests for both within-subject and leave-one-subject-out classification accuracy, enabling a comprehensive evaluation of the model’s performance across different scenarios. Additionally, for each test instance, we extracted embeddings of N-1 unrelated samples from the test set as inputs. This means, apart from the entire test set, the model evaluated by N-Way accuracy (where N=2, 4, 10 in our experiments) on the test set. We report these results in Appendix 0.G.
@@
Similar to the classification task, in the retrieval task, the objective is to retrieve the Top-K images most related to a given stimulus image via its corresponding EEG signal. This implies that by merely changing the text embeddings of image labels used during prediction, to image embeddings, we can transition the task from classification to image retrieval. The performance in retrieval tasks is superior to classification tasks as using image embeddings in training. We conducted a detailed comparison of different methods in terms of their Top-K retrieval accuracy in image retrieval tasks. Given that contrastive learning is known to be sensitive to batch size, we also compared the performance improvement of different methods under varying batch sizes (batch size=16, 1024) (Appendix 0.G).
@@
The generation task presents more challenges than the other tasks. In this phase, EEG representations are fed into a two-stage generator. For each image condition in the test set, we generate 1 images from 10 subjects based on the corresponding EEG signal. Subsequently, image retrieval is performed for each generated image. The Top-1 and Top-5 accuracies are calculated. Since we utilized CLIP for alignment and did not extract lower-level visual features from the EEG, this metric is particularly useful. It helps in evaluating the semantic alignment between the generated images and their original counterparts.
0.B.2 Computing methods implementation
In the upstream EEG encoder part, we compared various methods. For the B.D. method  [4], we replicated the network structure as described in the original work, with the difference being in the shape of the input data, due to the original study’s focus on MEG. It is worth mentioning that we used the retention test method in the testing process, so the strategy of joint subject training is not suitable for our task. We modify its subject-wise layer as an Input layer for modeling the time dimension. Regarding the  [46], we also reproduced the EEG encoder as described in their paper, using spatiotemporal convolution modules and EEG projection modules with the same convolution parameters. To ensure fairness, we did not use the same hyperparameters as in the original paper. Instead, we chose settings that yielded excellent results upon reproduction. Across all methods, we used identical hyperparameters, apart from the network structures. These included batch size, optimizer, initial learning rate, and temperature parameters.
0.B.3 Architecture details
Table 1: Brain module configuration for use with a target latent of size 1024
@@ 0.B.3 Architecture details
Spatial attention block (N, C, T) (N, C, T) 553,078
Input layer (N, C, T) (N, C, T) 62,750
Plug-and-Play Conv module (N, C, T) (N, H1, H2) 103,680
Temporal aggregation (N, H1, H2) (N, H1*H2) 0
@@
Module MLP AdaConv IL SA TOP-1 TOP-5
8.01 ±1.97 25.41±5.74
21.55±6.26 50.78±8.77
23.57±6.00 53.50±7.26
27.71±6.22 58.71±7.49
ATM-S 26.49±5.81 56.58±8.67
8.01 ±1.97 25.41±5.74
18.67±5.48 46.31±9.11
21.11±5.56 50.48±7.74
24.64±5.97 54.17±5.93
ATM-E 22.43±5.41 51.24±7.27
0.B.4 Training details
In the initial design of the EEG encoder module, we adopted two approaches to guide model predictions: text embedding and image embedding. Due to the differences in feature granularity, alignments favoring image embedding tend to perform better in image retrieval and classification tasks. During the training process using contrastive learning, we found that a batch size of 16 is a prudent choice across all models. A batch size of 1024 means that a sufficient number of samples are compared in a single training step, which requires the model to have a higher noise resistance capability. As observed in Tab. 5 and Tab. 6, larger batch sizes have the potential to achieve better performance and definite training efficiency gains across most methods. To improve the signal-to-noise ratio of EEG, we averaged 80 repeated data in the test set, a method similar to seeking Event-Related Potentials (ERP). To make full use of the training data volume, we did not average the 4 repetitions in the training set but instead fed all EEG data into the model for learning.
@@ 0.B.4 Training details
Figure 10: Test accuracy during training. (a) Training of within-subject model. (b) Training of across-subject model. We compared 5 different EEG encoding models, including EEGNetv4, B.D., NICE, ATM-S and ATM-E.
Appendix 0.C Details of EEG guidance image generation
@@
Here, we provide a concise overview of the conditional diffusion model framework used in EEG-guided image generation, following the presentation of continuous-time diffusion models in [45, 22].
@@
The core of DM training is to learn a model $s_{\theta}(x;t)$ for the score function. This is typically achieved through denoising score matching (DSM), where $\epsilon_{\theta}$ is a learnable denoiser. The training process can be formulated as:
$\mathbb{E}_{(x_{0},c)\sim p_{\text{data}}(x_{0},c),(n_{t},t)\sim p(n_{t},t)}\left[\|\epsilon_{\theta}(x_{0}+n_{t};t,c)-x_{0}\|_{2}^{2}\right],$ (3)
where $n_{t}$ is Gaussian noise with variance $\sigma_{t}^{2}$, and $c$ represents a condition.
0.C.1 Stage I - EEG-Conditioned Diffusion
@@
0.C.2 Stage II - CLIP-Embedded Image Synthesis
@@
Figure 11: Comparison between one-stage and two-stage EEG guidance image generation. (a) We present the images that subjects seen (Seen), the generated images directly using EEG embeddings (One-stage), and the generated images from image embeddings obtained by the prior diffusion (Two-stage). These results indicate that the strategy of our two-stage generation can better reconstruct the seen visual stimuli. (b) We employed ATM-S to compare the generated images with the original images in a retrieval task. Our result indicates that the images generated in two stages significantly enhance the performance of the original model on the retrieval task.
In the second stage of our EEG-based image generation approach, the CLIP embedding $z_{I}$ derived from the EEG-conditioned diffusion acts as the precursor for synthesizing visual objects $I$ based on $z_{I}$. This is achieved by harnessing the synergies of advanced pre-trained models, namely SDXL and IP-Adapter [33, 59], facilitating the creation of high-caliber images.
@@
To augment the efficiency of our framework, we additionally explore the SDXL-Turbo [36], a refined iteration of SDXL optimized for swift image synthesis. This variant proves especially beneficial in scenarios demanding quick generation of high-fidelity visuals.
@@
Appendix 0.D Performance comparison
@@
Our study uses various metrics to evaluate how well we can recreate visual stimuli from brain data (EEG, MEG, fMRI) (Table 1 in main text). These metrics include SSIM (structural similarity index metric)[51], SwAV (SwAV-ResNet50, refer to average correlation distance)[5], and two-way identification using neural networks (AlexNet(2/5), Inception, CLIP. Here AlexNet(2/5) the 2nd and 5th feature layers of AlexNet) for both low-level and high-level image features. Here two-way identification can be seem as a two-way retrieval task described in Appendix  0.B.1. In Tab. 3, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM. Tab. 4 shows the decoding performance of different data sets (fMRI, MEG, EEG) on visual stimulus tasks, and we even achieved the same or better performance than fMRI and MEG. Our results suggest that a suitable upstream neural representation plays a decisive role in the downstream task.
Table 3: Quantitative assessments of the reconstruction quality for EEG, MEG, and fMRI. For detailed explanations of the metrics.
Low-level High-level
Dataset $\uparrow$ SSIM $\uparrow$ AlexNet(2) $\uparrow$ AlexNet(5) $\uparrow$ Inception $\uparrow$ CLIP $\uparrow$ SwAV $\downarrow$
NSD (B.D.) [4] 0.366 0.962 0.977 0.910 0.917 0.410
NSD (Brain-Diffuser) [31] 0.356 0.942 0.962 0.872 0.915 0.423
NSD (MindEye) [39] 0.308 0.917 0.974 0.936 0.942 0.369
THINGS-MEG (B.D.) [4] 0.327 0.695 0.753 0.593 0.700 0.630
THINGS-MEG (Ours) 0.340 0.613 0.672 0.619 0.603 0.651
THINGS-EEG (Ours) 0.345 0.776 0.866 0.734 0.786 0.582
@@
GOD-Wiki (fMRI) CADA-VAE (V&T)[38] 10.02 40.37 - - - -
MVAE (V&T) [56] 10.04 39.60 - - - -
MMVAE (V&T) [43] 11.68 43.29 - - - -
MoPoE-VAE (V&T) [47] 12.90 51.78 - - - -
BraVL (V&T) [11] 13.99 53.13 - - - -
@@
Appendix 0.E Representational analysis
@@
Appendix 0.F Additional images results
0.F.1 Additional retrieval results

### v6

@@ B.1 Evaluation metric implementation
As CLIP has been designed to align text and image modalities, we also leverage its text encoder for EEG classification using the text embeddings of categories. This approach utilizes CLIP’s text encoding capabilities to facilitate EEG classification. We conducted zero-shot classification tests on the THINGS-EEG dataset. We employed Top-K accuracy as a metric for performance evaluation. Specifically, we assessed performance based on the Top-k (where k=1, 5) predictions. We conducted tests for both within-subject and leave-one-subject-out classification accuracy, enabling a comprehensive evaluation of the model’s performance across different scenarios. Additionally, for each test instance, we extracted embeddings of N-1 unrelated samples from the test set as inputs. This means, apart from the entire test set, the model evaluated by N-Way accuracy (where N=2, 4, 10 in our experiments) on the test set. We report these results in Appendix H.
@@
Similar to the classification task, in the retrieval task, the objective is to retrieve the Top-K images most related to a given stimulus image via its corresponding EEG signal. This implies that by changing the text embeddings of image labels to image embeddings, we can transition the task from classification to image retrieval. Given that contrastive learning is known to be sensitive to batch size, we also compared the performance improvement of different methods under varying batch sizes (batch size=16, 1024) (Appendix H).
@@
The generation task presents more challenges than the other tasks. For each image condition in the test set, we generate 10 different images from 10 subjects based on the corresponding EEG. Subsequently, image retrieval is performed for each generated image. The Top-1 and Top-5 accuracies are calculated. It helps in evaluating the semantic alignment between the generated images and their original counterparts.
B.2 Computing methods implementation
In the upstream EEG encoder part, we compared various methods. For the B.D. method  Benchetrit Y, (2023), we replicated the network structure as described in the original work, with the difference being in the shape of the input data due to the original study’s focus on MEG. It is worth mentioning that we used the leave one for subject method in the testing process so we modify its subject-wise layer as an linear layer for modeling the time dimension. To ensure fairness, we did not use the same hyperparameters as in the original paper. Instead, we chose settings yielded excellent results upon reproduction. Across all methods, we used identical hyperparameters, apart from the network structures. These included batch size, optimizer, initial learning rate, and temperature parameters.
B.3 Architecture details
Table 2: Brain module configuration
@@ B.3 Architecture details
Channel-wise attention layer (N, C, T) (N, C, D) 553,078
Temporal-Spatial Conv module (N, C, D) (N, H1, H2) 103,680
Temporal-Spatial aggregation (N, H1, H2) (N, H1*H2) 0
@@
Module MLP TSConv CAL TOP-1 TOP-5
8.11±1.74 26.83±4.78
21.65±6.22 51.34±9.83
ATM-S 23.73±7.62 52.71 ±9.71
28.64±6.39 58.47±8.97
8.11±1.74 26.83±4.78
17.95±5.95 43.10±8.63
ATM-E 23.73±7.62 52.71 ±9.71
24.92±6.00 54.78±8.49
B.4 Model configuration
To validate the efficacy of our EEG encoder, we experimented with a variety of empirical setups aimed at optimizing the model’s efficiency. we leverage joint subject training to adapt to new subjects. Once a model is trained, it can be used for both reasoning about known subjects (subject-specific tokens) and reasoning about unknown subjects (shared tokens). In the context of channel-wise attention layer, we explored four distinct approaches for enhancing downstream retrieval capabilities: leveraging subject-specific tokens for retrieval, averaging all tokens for a more generalized retrieval, flattening the entire token set for direct retrieval, and preserving token dimensions to feed into the Temporal-Spatial convolution module for feature integration. Notably, the strategy of preserving dimensions and using the Temporal-Spatial convolution module emerged as the most effective.
In our quest to optimize token embeddings, we experimented with a variety of approaches, including using $1\times 1$ convolution and linear layers. Under our framework, using linear layers performs better than convolutions. Moving beyond, we delved into the efficacy of diverse Feed Forward Networks (FFNs) within the Transformer encoder layer. Our findings indicated that an FFN tailored to the temporal dimension emerged as the superior option. We also conducted a comparative analysis of various positional encoding techniques. Interestingly, for Temporal-Spatial convolution utilized in retrieval tasks, the significance of positional encoding diminished.
In a further exploration, we examined the impact of deploying convolutions at various stages and even contemplated the complete removal of the convolution module. It was discovered that situating convolutions post the Transformer encoder layer yielded the most favorable outcomes. Conversely, a shift to a MLP or the removal of the convolution module led to a notable degradation in performance. Our assumption is though the convolution’s inherent translational invariance is compromised when the context of the time dimension is disrupted, the efficiency of parameters it possesses may confer a resistance to overfitting, thereby maintaining its effectiveness.
Table 4: Impact of each module on the result in different configurations.
Module Config Top-1 (std) Top-5 (std)
Channel-wise attention layer w/ mean token 7.29 (3.01) 23.11 (6.62)
w/ flatten token 14.85 (5.46) 37.56 (8.92)
w/ keep dim 28.64 (6.39) 58.47 (8.97)
Token embedding w/ conv1d 24.81 (7.29) 55.68 (9.08)
w/ linear 28.64 (6.39) 58.47 (8.97)
Feed Forward Network w/ temporal dim 28.64 (6.39) 58.47 (8.97)
w/ spatial dim 19.92 (7.74) 47.59 (11.5)
w/o 26.45 (7.73) 57.00 (9.95)
Position encoding w/ sinusoidal 27.96 (6.54) 58.16 (8.44)
w/ learnable absolute 26.66 (6.56) 56.95 (7.95)
w/o 28.64 (6.39) 58.47 (8.97)
Temporal spatial convolution w/ pre 14.23 (4.20) 37.44 (6.06)
w/ post 28.64 (6.39) 58.47 (8.97)
w/ both 25.37 (5.16) 57.07 (6.13)
w/o 23.73 (7.62) 52.71 (9.71)
B.5 Training details
In our EEG projector module, we integrated two distinct strategies for steering model predictions: text embedding and image embedding. Given the variance in feature granularity, we observed that alignments that prioritize image embedding excel in tasks of image retrieval and classification. Throughout the training phase, our experiments revealed that a batch size of 16 is a judicious selection for all models. Conversely, a batch size of 1024, which implies a substantial number of samples are processed in each training iteration, necessitates the model to exhibit a heightened capacity for noise resistance. In order to enhance the signal-to-noise ratio within EEG data, we implemented an averaging technique on 80 repeated instances within the test set. This approach mirrors the methodology employed in identifying Event-Related Potentials (ERP). To maximize the utilization of the available training data, we refrained from averaging the 4 repetitions in the training set. Instead, we opted to input the complete set of EEG data into the model, thereby facilitating comprehensive learning.
@@ B.5 Training details
Figure 9: Test accuracy during each training epoch. (a) Training of different within-subject models. (b) Training of different across-subject models. We compared 7 different EEG encoding models, including EEGconformer, MLP, EEGNetv4, B.D., NICE, ATM-S (Ours) and ATM-E (Ours).
Appendix C Details of EEG guidance image generation
@@
Here, we provide a concise overview of the conditional diffusion model framework used in EEG-guided image generation, following the presentation of continuous-time diffusion models in Song et al., (2020); Karras et al., (2022).
@@
The core of DM training is to learn a model for the score function. This is typically achieved through denoising score matching (DSM), where $\epsilon_{\theta}$ is a learnable denoiser. The training process can be formulated as:
$\mathbb{E}_{(x_{0},c)\sim p_{\text{data}}(x_{0},c),(n_{t},t)\sim p(n_{t},t)}\left[\|\epsilon_{\theta}(x_{0}+\sigma_{t}\epsilon;t,c)-\epsilon\|_{2}^{2}\right],$ (3)
where $\epsilon$ is Gaussian noise with variance $\sigma_{t}^{2}$, and $c$ represents a condition.
C.1 Stage I - EEG-Conditioned Diffusion
@@
C.2 Stage II - CLIP-Embedded Image Synthesis
@@
Figure 10: Comparison between one-stage and two-stage EEG guidance image reconstruction. (a) We present the images that subjects seen (Seen), our reconstructed images directly using EEG embeddings (One-stage), and the reconstructed images from low level and high level image embeddings obtained by the prior diffusion (Two-stage). These results indicate that the strategy of our two-stage generation can better reconstruct the seen visual stimulus. (b) We employed ATM-S to compare the generated images with the original images in a retrieval task. Our result indicates that the images generated in two stages significantly enhance the performance of the original model on the retrieval task.
Table 5: Latent VAE retrieval performance
Condition Top-1 (%) Top-5 (%)
ideal chance 0.5 2.5
ATM-S (Ours) 10.14 29.55
In the second stage of our EEG-based image generation approach, the CLIP embedding $z_{I}$ derived from the EEG-conditioned diffusion acts as the precursor for synthesizing visual objects $I$ based on $z_{I}$. This is achieved by harnessing the synergies of advanced pre-trained models, namely SDXL and IP-Adapter Podell et al., (2023); Ye et al., (2023), facilitating the creation of high-caliber images.
@@
To augment the efficiency of our framework, we additionally explore the SDXL-Turbo Sauer et al., (2023), a refined iteration of SDXL optimized for swift image synthesis. This variant proves especially beneficial in scenarios demanding quick generation of high-fidelity visuals.
@@
C.3 Low-level pipeline
Compared with pure vision pre-training models such as (ViT, ResNet, DINO, etc.), the CLIP model lacks low-level visual features. Therefore, in order to make up for this shortcoming, our framework introduces a low-level visual reconstruction pipeline. We hope to restore basic such as contour, posture, orientation and other pixel-level information from EEG by aligning with the latent of VAE.
Past work Takagi and Nishimoto, (2023) has found that in the denoising stage of the early diffusion model, $z$ signals (corresponding to the VAE latent in our framework) dominated prediction of fMRI signals. And during the middle step of the denoising process, $z_{c}$ predicted activity within higher visual cortex much better than $z$. However, note that this is only an analysis based on decoding accuracy. These analyses do not have a strong neuroscience causal relationship. We still cannot conclude that the low-level features of neural data are modeled by VAE.
We trained the low-level pipe for 200 epochs, trying a latent mean squared error (MSE) loss, along with a contrastive learning loss, and a variational autoencoder (VAE) image reconstruction loss to align the $4\times 64\times 64$ EEG latents obtained from a projection layer and an upsampled CNN with the VAE latents. Nevertheless, reconstruction loss or contrastive learning loss performs worse than only applying the loss in latent space and also requires significantly more GPU memory. In addition, we found that using a low-level visual model for distillation learning in the low-level pipeline is not only unhelpful for VAE latent training, but also leads to overfitting. Similar conclusions were reached in MindEyeScotti et al., (2023). Our results suggested that the low-level zeroshot reconstruction in EEG is not stable enough and may mislead the model results. When using the low level pipeline, we usually set the inference steps of SDXL to $10$ (or SDXL-turbo to $4$) and the image-to-image denoising strength to $0.5$. We give several reconstruction examples in Fig. 11 to compare the impact before and after using low level pipeline. Moreover, low level alignment was validated through retrieval performance tests, as depicted in Tab. 5. Our findings indicate that retrieval using EEG latents can also achieve excellent performance. This further substantiates the feasibility of aligning the low-level consistency through VAE latents.
Figure 11: Example of our reconstructions for Subject 8 output from different pipelines. From left to right: reconstruction using only CLIP (i.e., Only CLIP), using only CLIP and the semantic pipeline (i.e., CLIP $+$ Semantic), using only the low level pipeline (i.e., Only low level), using only CLIP and the low level pipeline (i.e., CLIP $+$ low level), using joint CLIP, low level, and semantic pipelines.
C.4 Semantic-level pipeline
In addition to using EEG latent and low-level pipelines during reconstruction, we also add a corresponding semantic level pipeline guided by text captions during the image reconstruction. We input the $1\times 1024$ EEG features output by prior Diffusion into the trained image projector to obtain $256\times 1024$ image features. Using the GIT modelWang et al., (2022), we can directly generate a caption from the latent features of the image. IP-Adapter accepts such a caption as a text prompt to guide the semantic level reconstruction of the image. It should be noted that due to the difficulty of the zeroshot task itself and the low dimensionality of the EEG features, the caption generated from the latent may be unstable, thereby interfering with the original correct EEG semantics. Considering that the image features extracted by the CLIP model itself are already high-level visual features and do not require the introduction of more semantic information, this framework retains the entry of the text prompt, and the reconstructed image presented does not force the use of the semantic level pipeline.
Appendix D Performance comparison
@@
Our study uses various metrics to evaluate how well we can recreate visual stimulus from brain data (EEG, MEG, fMRI) (Tab. 1 and Tab. 7). These metrics include PixCorr (pixelwise correlation, between ground truth and reconstructions), SSIM (structural similarity index metric)Wang et al., (2004), SwAV (SwAV-ResNet50, refer to average correlation distance)Caron et al., (2020), and two-way identification using neural networks (AlexNet(2/5), Inception, CLIP. Here AlexNet(2/5) the 2nd and 5th feature layers of AlexNet) for both low-level and high-level image features. Here two-way identification can be seem as a two-way retrieval task described in Ozcelik and VanRullen, (2023). In Tab. 1, our results showed that on the THINGS dataset, we could achieve performance over MEG on EEG reconstruction using ATM. Tab. 6 shows the decoding performance of different data sets (fMRI, MEG, EEG) on visual stimulus tasks, and we even achieved the same or better performance than fMRI and MEG. Our results suggest that a suitable neural representation plays a decisive role in the downstream task.
@@
GOD-Wiki (fMRI) CADA-VAE (V&T)Schonfeld et al., (2019) 10.02 40.37 - - - -
MVAE (V&T) Wu and Goodman, (2018) 10.04 39.60 - - - -
MMVAE (V&T) Shi et al., (2019) 11.68 43.29 - - - -
MoPoE-VAE (V&T) Sutter et al., (2020) 12.90 51.78 - - - -
BraVL (V&T) Du et al., (2023) 13.99 53.13 - - - -
@@
Table 7: Quantitative comparison results of image reconstruction in Subject 8 via our framework using different encoders.
Low-level High-level
Dataset $\uparrow$ PixCorr $\uparrow$ SSIM $\uparrow$ AlexNet(2) $\uparrow$ AlexNet(5) $\uparrow$ Inception $\uparrow$ CLIP $\uparrow$ SwAV $\downarrow$
NSD-fMRI Benchetrit Y, (2023) 0.305 0.366 0.962 0.977 0.910 0.917 0.410
THINGS-MEG Benchetrit Y, (2023) 0.058 0.327 0.695 0.753 0.593 0.700 0.630
THINGS-MEG (averaged) Benchetrit Y, (2023) 0.090 0.336 0.736 0.826 0.671 0.767 0.584
THINGS-MEG (Ours) 0.104 0.340 0.613 0.672 0.619 0.603 0.651
[…]
THINGS-EEG (EEGNetV4)Lawhern et al., (2018) 0.140 0.302 0.767 0.840 0.713 0.773 0.581
THINGS-EEG (Ours) 0.160 0.345 0.776 0.866 0.734 0.786 0.582
Appendix E Representational analysis
@@
Appendix F Concept analysis
We have adopted the concept embedding encoder proposed by Wei et al. Wei et al., (2024), which encodes the clip embedding of the original image into a 42-dimensional vector, with each dimension representing a distinct concept. Utilizing the ATM for direct projection on the EEG data of 200 categories from the test set, we obtained EEG embeddings that were then fed into the concept encoder with frozen weights, yielding 200 concept embeddings, as depicted in Fig. 13. An analysis of representational similarity at the concept level indicates that the EEG embeddings derived from our EEG projector effectively align with the conceptual space. This ensures semantic consistency at a high level of alignment, providing compelling evidence for the reconstruction of images from EEG data.
Figure 13: Visualization of the conceptual representation analysis. (a) Conceptual representations were obtained from eeg embeddings using concept encoder. (b) The similar matrix between EEG embeddings and real concept embeddings. (c) Concept embedding similarity matrix after cluster rearrangement (k=5).
Appendix G Additional images results
Figure 14: Examples of EEG-guided visual reconstruction. From top to bottom, we exhibit the best, median, and worst 12 generated images, respectively. We show the images subjects seen and the generated images by our two-stage image generator.
G.1 Additional retrieval results